# Identify As A Human Does: A Pathfinder of Next-Generation Anti-Cheat Framework for First-Person Shooter Games

### v1

@@ 6.8 Detecting Uncaught Cheaters in the Wild
User time (sec)1 415.72 178.12 36.89 3.97
System time (sec)2 69.32 436.61 27.12 13.53
Elapsed time (mm:ss)3 03:40 00:14 00:45 00:01
Percent of CPU this job got4 220% 4371% 140% 2860%
Maximum resident set size (GB)5 3.56 0.27 2.57 0.10
1User time: CPU time spent in user mode;
2
System time: CPU time spent in the kernel or system mode;
3
Elapsed time: Total real-world time taken to execute the job;
4
Percent of CPU this job got: How much of the available CPU resources were allocated to the process, this value is commonly greater than 100% when the process runs on multiple CPU cores;
5
Maximum resident set size: The peak amount of physical memory that the process used.
6.9 Hawk’s Overheads
Hawk is developed as an asynchronous, load-balancing system that analyzes post-game replay files during periods of low server activity. It operates independently of active game servers and can be deployed on a separate server. However, we still conduct overhead assessments for reference. To answer RQ.6, we utilize an Ubuntu server equipped with 128 cores AMD EPYC 7713P 64-Core Processor and 256 GB RAM to mimic a real dedicated game server [19]. By leveraging the pre-trained model in the artifact, Hawk requires only CPU computing resources to function gracefully. Table 7 indicates the maximum overhead statistics per match, with each match analyzed concurrently. In addition to the basic overheads, we also estimate the Hawk-to-server overheads in terms of a real deployment scenario. Our platform partner reports that at most 150,000 matches are played daily. Each server needs to run at most 573 matches a day. According to our platform’s report and the public server statistics [56], player activity is at its lowest, typically 1-5 AM local time. During this timeframe, at most 144 matches need to be analyzed per hour. Thus, according to the elapsed time mentioned in Table 7, at most it needs to process 12 matches concurrently to finish processing all matches within the timeframe per server. We further conduct concurrent processing evaluation, it occupies at most 17.95% memory and 35.45% CPU usage for processing 12 matches concurrently. Therefore, each server possesses more than sufficient resources for deploying Hawk, even if several matches are hosted at the same time.
7 Discussions
Generalizability. Narrowed generalizability is a meta-problem for the anti-cheat field. Since the cheat is game-specified, the anti-cheat is hard to generalize for all games. Although we successfully implemented Hawk on one of the leading tactical shooter games [64] (a subgenre of FPS games like CS:GO), like other prior works [9, 69, 70, 2, 24, 39] discussed in Section 8, it requires further adaptation for other FPS subgenres with distinct gameplay. Hawk’s generalizability concerns on the core observations, methodologies, and workflow rather than specific features. The features presented are theoretically applicable to tactical shooters in FPS games like the CS series, CrossFire, and others. However, adaptation is needed for other subgenres, where certain elements like flashes and grenades may not apply. For instance, Hero Shooters (e.g., Overwatch, Valorant) where such features could be substituted with character-specific abilities, or Battle Royale (BR) games (e.g., Fortnite, PUBG) where elements like angles and velocities need substitution due to long time-to-kill. Since the replay files and labelings are significant for Hawk, also due to the dataset scarcity limitation, conducting further research on different subgenres to prove the generalizability of Hawk must require cooperation with the game vendors to obtain the data. We have initiated contact with several vendors, and one specializing in BR games has shown interest in cooperating. But, we have to leave specific implementation and evaluation for future work.
Manual involvement. Although Hawk has significantly increased the detection performance and efficiency, as mentioned in Section 3.2, Hawk-after-use manual verification is unavoidable. The presence of real false positives undeniably indicates the requirement for manual checks to avoid false bans. Nevertheless, with the mitigation of the limitations mentioned in Section 1, we still believe Hawk plays an important role in the FPS games’ server-side anti-cheat.
8 Related Works
Client-side detection. BotScreen [9], as the latest aimbot detector, utilizes TEE (Intel SGX) to store and run the detection model with aiming features. It leverages a self-setup server and hires 14 players to play in Deathmatch instead of the rank mode. Thus, BotScreen is limited to dataset scarcity, and its real-world functionality remains unclear, especially with more complex scenarios. BotScreen evaluates with a basic, open-source aimbot Osiris [37], which is game-specific and may not reflect the performance against more sophisticated cheats [44, 42]. VAC [12] can detect wallhack by using OpenGL and scanning such as DNS cache to verify if the suspect has purchased and cheated in a game for Counter-Strike [38, 12]. Warden [61] scans the system running programs, users’ browser history, and cookies to determine a user’s behavior for Blizzard’s games. EAC [22], BattlEye [3], and XIGNCODE3 [62] are commercial anti-cheats. They can achieve both cheat detection and prevention [53] for multiple platforms and different games after customization by monitoring and managing process privileges. Vanguard [30] is a kernel-level anti-cheat for VALORANT [51] that has full access and control over the system. The above client-side anti-cheats are limited to hardware constraints, security and privacy concerns, and client overheads. And reflect the limited generalizability problem for anti-cheat.
Client-side prevention. BlackMirror [49] applies SGX for anti-wallhack, providing the client GPU only visible entities based on local predictions. It only prevents without detecting and is limited to hardware constraints, client overheads, and targeting on a specific cheat type or subcategories. Invisibility Cloak [58] prevents computer-vision-based aimbots from performing object detection with the screen images but is ineffective against memory-based aimbots and wallhacks.
Server-side ML-based detection. Yeung et al. [69] initially proposed a Bayesian model to tune the probability threshold with basic statistics in 2006, but it is outdated for today’s advanced cheating techniques. Subsequent works [70, 2, 24, 39] used machine learning with basic statistics or limited features, which couldn’t monitor sophisticated behaviors or describe entire matches [9]. These earlier works are limited to naive features and workflow, unclear performance, and dataset scarcity. They require extra modification (specified to the game) on the game engine to collect and construct features in parallel with the game commencing. Therefore, it is also hard to generalize their method to other games.
9 Conclusion

### v2

@@ VI-I Overheads
User time (sec) 415.72 178.12 36.89 3.97
System time (sec) 69.32 436.61 27.12 13.53
Elapsed time (mm:ss) 03:40 00:14 00:45 00:01
Percent of CPU this job got 220% 4371% 140% 2860%
Maximum resident set size (GB) 3.56 0.27 2.57 0.10
TABLE XI: Performance under varying concurrency levels.
# Concurrent Request 10 20 50
Average time (sec) 384.512 365.267 377.084
Minimum time (sec) 380.343 357.570 368.070
Maximum time (sec) 393.472 391.536 386.522
Standard deviation 4.502 7.793 5.282
VII Discussions
Generalizability. Narrowed generalizability is a meta-problem in the anti-cheat field. Since cheats are game-specific, creating a universal anti-cheat system is challenging. Although we successfully implemented Hawk on a leading tactical shooter game [65] (a subgenre of FPS games like CS:GO), similar to prior works [6, 10, 11, 12, 13] discussed in Section VIII, it requires further adaptation for other FPS subgenres with distinct gameplay. The proposed features are applicable to tactical shooters with different engines in FPS games, e.g., CS2. However, adaptation is needed for other subgenres, where certain elements like flashes and grenades may not apply. We summarize two core principles for adapting Hawk to other subgenres. (a) Generalized anti-cheat perspectives transform from three observations. The fundamental structure of transforming three key observations into three corresponding anti-cheat subsystems is a generalized framework applicable to any FPS subgenre. While the specific features or model architectures within each subsystem may need to be optimized for different games, the principle of using these three complementary perspectives to identify cheating behavior is universally necessary and applicable. (b) Adaptable feature extraction strategy. We propose that feature extraction for any FPS game that uses Hawk should be divided into two distinct categories. Core FPS features. These are fundamental features inherent to all FPS games and should be retained across all subgenres. For example, this includes data like player position, distance to target, and view direction (in RevPov); or various aiming speeds, angles, and precision/hit rates (in RevStats). While generalizing Hawk to other subgenres, these features are applicable without any modification. Subgenre-specific features. These features are designed around the unique mechanics that define a specific subgenre. The design of subgenre-specific features, as well as the classification rules in ExSPC, should be led by senior anti-cheat developers with deep expertise in the specific game. This ensures the system is tailored to detect cheats that exploit that subgenre’s unique mechanics. For instance, in a Hero Shooter (e.g., Overwatch, Valorant), the ability to use unique character skills is a key differentiator. To adapt Hawk to other subgenres, one would: i. augment RevPov with basic temporal information related to skill activation and targeting; ii. extend RevStats to record metrics like skill accuracy, the mean and variance of skill usage timing, and their effectiveness.
Manual involvement. Although Hawk has significantly increased the detection performance and efficiency, as mentioned in Section III-B, Hawk-after-use manual verification is unavoidable. However, platform statistics indicate their anti-cheat recall (47%) is significantly lower than Hawk’s (79% on average) under recall-first mode. Although Hawk ’s false FPR is around 25%, the high recall outweighs the FPR increase (about 15%). A higher FNR indicates more cheaters undetected, leading to user attrition with greater losses than those caused by higher FPR [66]. Hawk also offers Balance mode whose FPR is lower and recall is higher than the current in-use anti-cheat (Section VI-F and footnote 4). Under this mode, Hawk outperforms current anti-cheat and obtains lower manual involvement, due to lower FPR. Note that for any anti-cheat system, all results must undergo manual review and confirmation before any bans are issued (i.e., the GM verification stage). To provide an intuitive demonstration of Hawk’s performance, unless otherwise specified, all results presented for this system are shown before manual review. Therefore, all positive samples mentioned in the evaluation require manual review before any bans can be enforced in actual operation. No player will be mistakenly banned or inconvenienced before the manual revision. Regarding the trade-off between recall and FPR, we have demonstrated the system’s performance under different anti-cheating requirements in Table VIII. It can be observed that for every approximately 5% increase in recall, FPR increases by 10% and 20.00% and 20.84% increase of the manpower required on average in terms of the aimbot and the wallhack, respectively. Therefore, if human resources are relatively scarce, the default or balanced mode of Hawk should be used.
Dataset imbalance and result significance. To statistically validate the performance stability of our models, we employed bootstrap resampling [67] with 5,000 iterations to estimate 95% confidence intervals for all evaluation metrics in Table XII. The bootstrap analysis confirmed the statistical significance and robustness of our models under class-imbalanced conditions. All performance metrics showed narrow 95% confidence intervals. This demonstrates that both aimbot and wallhack detection performances are not only practically meaningful but also statistically reliable, validating the effectiveness of our approach even with inherent dataset imbalances in the context of game anti-cheating.
TABLE XII: Performance metrics with 95% confidence intervals from bootstrap analysis (5,000 resamples).
Dataset Metric Original Bootstrap_Mean Std_Dev Std_Error CI CI_Width
Aimbot Accuracy 0.716149 0.716142 0.00467 0.000066 [0.706972, 0.725323] 0.018351
Recall 0.726244 0.726073 0.014762 0.000209 [0.697646, 0.755435] 0.057789
NPV 0.960191 0.960179 0.00249 0.000035 [0.955263, 0.964983] 0.00972
AUC-ROC 0.72065 0.720535 0.007924 0.000112 [0.704752, 0.736172] 0.03142
FPR 0.284944 0.284831 0.004967 0.00007 [0.274992, 0.294452] 0.01946
Wallhack Accuracy 0.756375 0.756422 0.004182 0.000059 [0.748394, 0.764651] 0.016257
Recall 0.847805 0.847965 0.011178 0.000158 [0.826003, 0.869812] 0.043808
NPV 0.977897 0.977882 0.001741 0.000025 [0.97437, 0.981204] 0.006834
AUC-ROC 0.797024 0.797097 0.006064 0.000086 [0.785037, 0.80896] 0.023923
FPR 0.253757 0.253778 0.004606 0.000065 [0.244804, 0.262858] 0.018054
VIII Related Works
Server-side anti-cheats. Yeung et al. [10] initially proposed a Bayesian model to tune the probability threshold with naive features (e.g., whether the player is moving, the distance between the attacker and victim) in 2006, but its performance lacks real-world validation and the features are outdated for today’s advanced cheating techniques and sophistication. Subsequent works [11, 12, 13, 31, 30] employed statistical tests with basic features (e.g., time-on-target) to examine data distribution difference [11], set fixed threshold with limited features (e.g., extra shots count on an invisible bait) [13, 30], or used one machine learning model (e.g., SVM) with basic non-temporal features (e.g., whether the target’s position changed) [12], which couldn’t monitor sophisticated behaviors or describe entire matches [6], or ineffective to certain type of cheat. Due to the scarcity of real-world datasets, these server-side studies relied on private servers and simulated data for evaluation, with the largest dataset containing only 47 cheaters [13]. Such small-scale datasets are insufficient to reflect real-world anti-cheat performance and possibly lead to inconsistent performance between the reported results and real-world performance. Additionally, works like [11, 12, 13, 31] require game engine modifications for concurrent data collection, necessitating significant adaptation efforts after each update and limiting generalizability to other games. Importantly, prior works leverage a single model to classify and could result in biased predictions on the complex real-world dataset as discussed in Section V-B. CUPID [68] is a matchmaking design to optimize the user’s gameplay experience. In terms of anti-cheat, it may be modified to arrange cheaters within the same match so that they would not appear in the normal matches. However, its applicability remains unknown since it does not target identifying cheaters.
Client-side anti-cheats. BotScreen [6], the latest aimbot detector, uses SGX to securely store and run its detection model with temporal aiming features. It constructs its dataset using a self-setup server with 14 players in Deathmatch (gameplay is different from Rank) and evaluates with the deprecated open-source aimbot Osiris [69]. This setup may not accurately reflect performance against more sophisticated cheats or common cheating scenarios [70, 71], leaving its real-world functionality unclear. It is also troubled by the deprecated TEE constraint [14, 15]. Invisibility Cloak [8] and BlackMirror [9] only prevent without detecting and are limited to hardware constraints or client overheads, and targeting on a specific cheat subcategory. Commercial anti-cheats like VAC [18], Vanguard [35], EAC [16] and BattlEye [17] offer both detection and prevention [72] for various games through customization. They monitor and manage process privileges, scan DNS caches and cookies for browsing history and cheat purchase records [73, 18], and may use kernel-level drivers [33]. However, these solutions are limited by security and privacy concerns mentioned in Section I. The common approach of creating blacklists for known cheats and issuing bans based on detected malware signatures [38, 33, 6, 13] is ineffective against new or updated cheats with different signatures [6, 13], lacking robustness to cheat evolutions. Most cheaters avoid using legacy cheats as paid cheats often offer auto-updates to evade bans, leading to low recall in industry anti-cheats, which are always one step behind. These anti-cheats are integrated into games before release and must run concurrently with gameplay [33], causing client overheads and limited generalizability. They also combine multiple methods for cross-checking, extending the banning cycle [33]. Importantly, these commercial anti-cheats are proprietary and closed-source [33], making it infeasible for researchers to conduct comparison studies without collaborating with game vendors. VESPA [74] combines computer vision techniques and human inspection to capture the user’s screen and detect the wallhack’s UI overlays and thereby identify cheaters. This may raise privacy concerns and can be easily bypassed by managing process privileges or disabling the visible overlays. Therefore, it is less practical for real-world deployment. FCDGame [75] leverages few-shot learning to learn highly repetitive gesture trajectory patterns on the mobile phone screen to identify cheaters. This pattern targets the bot identification in MMORPGs rather than FPS games. Because the former only requires fixed navigation and simple gesture operations to achieve repetitive tasks (e.g., collect mines), but the latter requires real-time and complex inputs to adjust aiming and perform shooting. The FPS trajectory patterns are more complex and irregular. Therefore, the screen gesture trajectory feature may be less functional in FPS games.
IX Conclusion